if the number of ways of selecting n cards out of unlimited number of cards bearing the numberr 0,9,3, so that they cannot be used to write the number 903 is 93, then n is equal to ?
 

Grade:12th pass

1 Answers

ajay
7 years ago
formula to find the term is
3(2^n-1)=93
2^n-1-=31
2^n=32
2^n=2^5
n=5
so you will get required answer 5